OpenStudy (anonymous):

A car travels at 20 meters per second. What is its rate in kilometers per hour?

Parth (parthkohli):

I don't know if you know this, but we multiply it by 18/5 to convert into km/h

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {20 } \cdot {18 \over 5}}\)

Parth (parthkohli):

@shubham.bagrecha I believe something is wrong with that

Parth (parthkohli):

\( \color{Black}{\Rightarrow 4 \cdot 18}\) after cancelling ^

OpenStudy (anonymous):

Metre/Second is given.. To convert Metre into Kilometre, we divide it by 1000 And to convert Second into Hour, we divide it by 3600 So, Meter/Second = (1/1000)/(1/3600) Kilometer/Hour So, Meter/Second = 3600/1000 = 36/10 = 18/5 Kilometer/Second So, simply you can multiply the given with 18/5 to get the desired result.. Parthkohli is saying absolutely right...

OpenStudy (anonymous):